in a viral pool test it is known that in a group of 5 people, exactly one will test positive. if they are tested one by one in a random order for confirmation what is the probability the only two test are needed


1
Expert's answer
2021-06-07T04:56:24-0400

The first occurrence of success requires 2 trials


"P=(1-\dfrac{1}{5})(\dfrac{1}{5-1})=\dfrac{1}{5}"